OpenStudy (anonymous):

Find a pair of factors for each number by using the difference of two squares. a. 45 b. 77 c. 112

OpenStudy (anonymous):

a. 49 -4 = 7² - 2² = (7+2)(7-2) ??? am i on the right track with these???

OpenStudy (anonymous):

b. 81 -4 = 9^2 - 2^2 = (9+2)(9-2) c. 121 -9 = 11^2 - 3^ = (11+3)(11-3)

OpenStudy (anonymous):

Let's start with a. Using difference of two squares, note that 45=49−4=72−22This last expression can be factored easily as 72−22=(7−2)⋅(7+2) Hence, your factorization is 9⋅5

OpenStudy (anonymous):

For b, we want to write 77=81−4=92−22 There's an easy way to factorize a difference of two squares. If you have a differnece of two squares a2−b2 then this factors into a2−b2=(a−b)⋅(a+b) Does this help? So you have an equation of the form a2−b2 where a=9 and b=2. Thus, a2−b2=(a−b)⋅(a+b) can be replaced by 92−22=(9−2)⋅(9+2) Which then simplifies to 7⋅11 Let's move on to c. We want to factor 112. Notice that 112=121−9=112−32 Using the same principles as above,

OpenStudy (anonymous):

c. 80, 32 ???

OpenStudy (anonymous):

c. 8, 14 ????

OpenStudy (anonymous):

i'm so confused!

OpenStudy (anonymous):

Wait , let me break this down to you. a. 45 = (7-2)(7+2) b. 77 = (9-2)(9+2) c. 112 = (11-3)(11+3) so therefore, if you go with c , it may likely be your answer , lol im kinda confused here as well